OpenStudy (vortish):

if possible evaluate the expression for the given value of the variable 2/x^2-4x-5 x =1

OpenStudy (aripotta):

2/1^2-4(1)-5

OpenStudy (aripotta):

so... 2/1-4-5 2-4-5 -7?

OpenStudy (vortish):

1-4-5 x=-9 ?

OpenStudy (aripotta):

i though x = 1?

OpenStudy (aripotta):

thought*

OpenStudy (vortish):

it is x =1

OpenStudy (aripotta):

so the answer's -7, right? or did i do something wrong? :/

OpenStudy (vortish):

not the right answer

OpenStudy (aripotta):

oh. herm.

OpenStudy (aripotta):

is it like 2 over everything else?

OpenStudy (aripotta):

 like this?

OpenStudy (aripotta):

or wut

OpenStudy (vortish):

have a new expression because I did not get it right 10/x^2=5x+4 x=1

OpenStudy (aripotta):

that's an equation

OpenStudy (vortish):

ok equation

OpenStudy (aripotta):

but. what exactly are we doing with this? we can't solve an equation. the variable's already solved for

OpenStudy (vortish):

if possible evauluate the expression for the given value of the variable

OpenStudy (aripotta):

but that's an equationnnnnn

OpenStudy (vortish):

just what it says ari

OpenStudy (aripotta):

but there's nothing we can do lol

OpenStudy (aripotta):

did you type it right?

OpenStudy (vortish):

if possible evaulate the expression for the given value of the variable for x =1 10/x^2-5x+4 simplyfi your answer

OpenStudy (aripotta):

thar she blows

OpenStudy (aripotta):

 is it like this?